MARS DESERT RESEARCH STATION

Sol Summary – March 21st

MDRS Daily Summary Report for SOL 10

Person filling out Report: Karolina Zawieska, Green Hab Officer/Human-Robot Interaction Officer

Summary Title: White Canyon Exploration and outreach activities

Mission Status: Work in progress

Sol Activity Summary: EVA, documentation activities, cleaning

Look Ahead Plan: Continue EVA and documentation activities, solve the Internet problem

Anomalies in work: Unstable Internet connection

Weather: Cloudy in the morning, windy but sunny in the afternoon

Crew Physical Status: One of the crew members suffering from mild knee pain, everyone else feels good

Reports to be filled:

-Operations Report
-Commander Report
-EVA Report
-EVA Request

Sol Summary – March 18th

MDRS Daily Summary Report for SOL 7

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title: Explore and relax

Mission Status: recharged, thinking about everything except problems

Sol Activity Summary: human-robot interactions on EVA, relax, maintenance

Look Ahead Plan: EVA,  media and outreach activities

Anomalies in work: due to accidental software update, we’ve lost almost 30% of our daily internet data allowance.

Weather: sunny, small clouds

Crew Physical Status: everyone is alive

Reports to be file:
-Operations report
-Commander report
-EVA Report
-EVA Request
-Science report (geology)
-Science report (engineering)
-Science report (Human-Robot Interaction)

Sol Summary – March 17th

MDRS Daily Summary Report for SOL 6

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title: It’s a good day to explore Mars

Mission Status: we can focus more on our field work instead of current hab maintenance

Sol Activity Summary: geo research during EVA, installing new pump, other work in hab with robot

Look Ahead Plan: EVA, mid-term summary session

Anomalies in work: our first day with every hab system working nearly properly

Weather: sunny, no strong wind

Crew Physical Status: one of crew members has mild cold sympthoms

Reports to be file:

-Operations report
-Commander report
-EVA Report
-EVA Request

Sol Summary – March 16th

MDRS Daily Summary Report for SOL 5

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title:  Team is finally complete

Mission Status: successful EVA, progress with current work

Sol Activity Summary: training for 5th crew member, EVA, maintenance, green hab work

Look Ahead Plan: further work with robot, EVA,

Anomalies in work: mysterious water leakage outside the hab getting even bigger, need to manually refill static tank without a pomp

Weather: sunny, winds in the evening

Crew Physical Status: everything is correct

Reports to be file:

-Operations report
-Commander report
-Green Hab Officer’s report
-EVA request

Sol Summary – March 13th

MDRS Daily Summary Report for SOL 2

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title: Last day of freedom

Mission Status: Final prepearation for SIM

Sol Activity Summary: reconnaissance outside the hab (and surrounding roads), fixing the tunnel, geological survey

Look Ahead Plan: first day in fully isolation (SIM)

Anomalies in work: solar panels still don’t work properly

Weather: sunny, calm, no strong wind

Crew Physical Status: minor injury while opening the can, but everyone is still alive and in good condition

Reports to be file:

-Operations report

-Commander report

-Mission Plan

-Bios and Photos

-Science report

Sol Summary – March 2nd

SOL SUMMARY REPORT SOL 18
SOL: 18

Person filling out Report: Louis MALLER, XO

Summary Title: The Artists

Mission Status: successful silent EVA in the morning, all systems go, work on experiments ongoing

Sol Activity Summary: EVA, science work on different experiments, sport

Look Ahead Plan: Tomorrow we will go for a last EVA, debrief the mission, and prepare to exit the Sim.

Anomalies in work: gave up on the 3D printer, clock in the living quarters not functional

Weather: sunny, no wind

Crew Physical Status: Injured crewmember conditions stable, one crewmember with a slight headache, rest of crew feeling well

EVA: non-verbal EVA for balloon debris recovery, AR glasses tests and seismometer recovery – success

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #17 report
– EVA #18 request
– Science Report

Sol Summary – March 1st

SOL SUMMARY REPORT SOL 17
SOL: 17

Person filling out Report: Louis MALLER, XO
Summary Title: Up!
Mission Status: successful EVA in the morning, all systems go, work on experiments ongoing
Sol Activity Summary: EVA, science work on different experiments, sport
Look Ahead Plan: Tomorrow we will gather the seismometer, then explore some more as we prepare to leave
Anomalies in work: gave up on the 3D printer, clock in the living quarters not functional, balloon untethered and seems to have flown away
Weather: very sunny, some mild winds
Crew Physical Status: Injured crewmember conditions stable, rest of crew feeling well
EVA: EVA for balloon deployment, AR glass tests and some around the Hab exploration

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #16 report
– EVA #17 request
– Astronomy Report
– Science Report

Sol Summary – February 28th

SOL SUMMARY REPORT SOL 16
SOL: 16

Person filling out Report: Louis MALLER, XO

Summary Title: Rose-tinted glasses

Mission Status: successful EVA in the afternoon, all systems go, work on experiments ongoing

Sol Activity Summary: sport, EVA, science work on different experiments

Look Ahead Plan: Hoping to have good weather for 24-hour balloon deployment tomorrow

Anomalies in work: 3D printer being troubleshooted, clock in the living quarters not functional

Weather: very sunny, some mild winds

Crew Physical Status: Injured crewmember conditions stable, rest of crew feeling well

EVA: EVA for engineering check and some pictures for emergency protocols report

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #15 report
– EVA #16 request
– Science Report

Sol Summary – February 27th

SOL SUMMARY REPORT SOL 15
SOL: 15
Person filling out Report: Louis MALLER, XO
Summary Title: Betelgeuse Broadcasting Corporation
Mission Status: successful EVA in the late morning, all systems go, work on experiments ongoing
Sol Activity Summary: brunch, EVA, science work on different experiments
Look Ahead Plan: Hoping to have good weather for 24-hour balloon deployment
Anomalies in work: 3D printer being troubleshooted, clock on the first floor not working correctly (little arm stuck between 6 and 7 O’clock)
Weather: mostly sunny, very strong gusts of winds.
Crew Physical Status: Injured crewmember feeling better, rest of crew feeling well
EVA: EVA with visiting journalists from the Betelgeuse Broadcasting Corporation, sent by Satellite 5, checked the seismometer, tested the alternative Martian navigation system, and explored moon- and mars- like regions.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #14 report
– EVA #15 request
– HSO report
– Astronomy Report
– Science Report

Sol Summary – February 26th

SOL SUMMARY REPORT SOL 14
SOL: 14
Person filling out Report: Louis MALLER, XO
Summary Title: Brunch on Mars
Mission Status: successful EVA in the afternoon, all systems go, work on experiments ongoing
Sol Activity Summary: brunch, EVA, science work on different experiments
Look Ahead Plan: Looking forwards to Brian Cox joining us tomorrow for a Sol on Mars.
Anomalies in work: 3D printer being troubleshooted
Weather: beautiful and sunny, breeze with gusts of winds.
Crew Physical Status: Injured crewmember feeling better, rest of crew feeling well
EVA: EVA in front of the Hab, in order to learn, test and refine security procedures, went quite well.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #13 report
– EVA #14 request
– HSO report
– Astronomy Report